# DCIM-AVSR : Efficient Audio-Visual Speech Recognition via Dual Conformer Interaction Module

### v1

@@ V Conclusion
In this paper, we propose a Dual Conformer Interaction Module (DCIM) to enhance multimodal interaction. Building on this module, we introduce an asymmetric DCIM-AVSR model, where the visual modality serves as the auxiliary mode and the audio modality as the primary mode. The model distributes the process of modal information fusion across each DCIM module. Additionally, we introduce a pre-training strategy that maximizes the benefits of the pre-trained model, reducing both learning and training costs while improving efficiency. Our experiments show that the DCIM-AVSR model achieves the same robustness and even better performance compared to conventional AVSR models.

### v5

@@ V Conclusion
In this paper, we introduce an asymmetric DCIM-AVSR model, where the visual modality serves as the auxiliary modality and the audio modality as the primary modality. We also propose a Dual Conformer Interaction Module (DCIM) to enhance multi-modal interaction. The model distributes the information fusion process across each DCIM module. Additionally, we introduce a pre-training strategy that maximizes the benefits of the pre-trained model, reducing learning and training costs while improving performance. Our experiments show that the DCIM-AVSR model achieves better performance compared to conventional AVSR models with great robustness.